# DoubleAgents: Interactive Simulations for Alignment in Agentic AI

### v2

@@ 9. Conclusion
We introduced DoubleAgents, a system for aligning agentic AI through interactive simulation, policy-grounded planning, and transparent oversight. In coordination tasks like speaker scheduling, alignment is not just about correctness; it’s about respecting user preferences, adapting to edge cases, and maintaining trust. By embedding simulation into the agent workflow, DoubleAgents helps users rehearse edge cases, refine their preferences, and gradually scale up delegation. Our lab and real-world deployments show that users increasingly trusted and customized the system, not because it was perfect, but because they could iteratively shape it. This work offers a practical path toward alignment in real-world agentic AI: one grounded not in static prompts, but in interaction, iteration, and co-adaptation.
10. Generative AI Usage Disclosure
We used generative AI tools (e.g., OpenAI’s ChatGPT) as support during several parts of this project. Specifically, we used them to help debug the system, draft non-critical implementation and UI code, write analysis scripts for logs and study data, and edit text for grammar and clarity during paper writing.
@@ Appendix A Seminar Planning: A Human Coordination Problem
Thus, experienced organizers adopt personal strategies to cope with such a demanding task. For example, some follow a rule of “wait two days, then follow up” to nudge busy people who may not read emails right away, while others wait a full week before reaching out again. Some prefer contacting close collaborators or junior researchers first, while others prioritize securing prominent professors or distant speakers early to reduce travel constraints. When conflicts arise between speakers, organizers might “give faculty priority and ask students to be more flexible,” carefully deciding whom to approach for alternative availability to minimize disruption. For non-responders or last-minute cancellations, they must judge when and how often to follow up before moving on to new options—and sometimes even decide when it is appropriate to “beg” busy invitees for additional availability. Because plans are constantly evolving, organizers become locked in a cognitively and time-intensive loop of tracking shifting states, replanning, and executing revised plans as new information arrives—yet they often cannot proceed with the next steps or finalize allocations until replies are received.
Since seminar scheduling requires nuanced, context-dependent decisions that reflect each organizer’s values and preferences, it is difficult to fully automate. An effective seminar scheduling tool should not attempt to handle the process independently; instead, it must learn from and adapt to the organizer, respecting social norms, relational dynamics, and the inherent unpredictability of responses.
Appendix B User Study Procedures and Data Analysis
@@
Appendix C DoubleAgents Prompts
C.1. Coordination Agent
C.1.1. Progress Summary
@@ C.1.4. Email Drafting
C.1.5. Implicit User Preference Learning
@@
C.2. Edge Case Detection
@@ C.3. Simulated Respondents
C.4. Free-form QA Chat
@@ C.5. Daily Summary
Appendix D Technical Evaluation Details
D.1. Policy Selection
@@ D.1.1. Experiment prompts
D.2. Edge Case Detection
D.2.1. Test suite construction
@@
D.2.2. Prompts to generate linguistic variations and test data

### v3

@@ 9. Conclusion
We present DoubleAgents, an interactive alignment tool for human coordination through user intervention, value-reflecting policies, rich state visualizations, and uncertainty flagging. Technical evaluations validate key system capabilities of DoubleAgents, including policy selection and edge-case detection. A two-session user study (N=10) shows users’ increased willingness to delegate both planning and execution, supported by policy-based alignment, interpretable visualizations, and explicit delegation boundaries. Three real-world deployments show that alignment learned through simulation transfers to live coordination while continuing to evolve in practice. Our findings frame alignment not as a one-time specification problem but as an iterative process, and position distributed cognition as a principled approach for enabling progressive delegation with retained user control.
@@ Appendix A Seminar Planning: A Human Coordination Problem
Thus, experienced organizers adopt personal strategies to cope with such a demanding task. For example, some follow a rule of “wait two days, then follow up” to nudge busy people who may not read emails right away, while others wait a full week before reaching out again. Some prefer contacting close collaborators or junior researchers first, while others prioritize securing prominent professors or distant speakers early to reduce travel constraints. When conflicts arise between speakers, organizers might “give faculty priority and ask students to be more flexible,” carefully deciding whom to approach for alternative availability to minimize disruption. For non-responders or last-minute cancellations, they must judge when and how often to follow up before moving on to new options—and sometimes even decide when it is appropriate to “beg” busy invitees for additional availability. Because plans are constantly evolving, organizers become locked in a cognitively and time-intensive loop of tracking shifting states, replanning, and executing revised plans as new information arrives—yet they often cannot proceed with the next steps or finalize allocations until replies are received. Since seminar scheduling requires nuanced, context-dependent decisions that reflect each organizer’s values and preferences, it is difficult to fully automate. An effective seminar scheduling tool should not attempt to handle the process independently; instead, it must learn from and adapt to the organizer, respecting social norms, relational dynamics, and the inherent unpredictability of responses.
Appendix B System Implementation Details
Figure 4. System architecture of DoubleAgents. The Policy Module supplies policies to the Coordination Agent within the Simulation Engine, which exchanges emails with Simulated Respondent Speakers to advance the task. When a response falls outside existing policy coverage, the Edge Case Detector escalates the situation to the user and feeds the resulting clarification back into the Policy Module as an updated policy. The Dashboard Visualization renders the evolving task state throughout this cycle.
We present DoubleAgents, a system for aligning agentic AI through interactive simulation. Built on the ReAct framework (Yao et al., 2023), DoubleAgents supports iterative planning and execution via a team of AI agents that collaborate with users on complex coordination tasks. DoubleAgents takes as input a user-defined goal (e.g., “Organize a seminar with four speakers and four slots”), persona descriptions for simulated respondent agents, the user’s existing email templates, and a set of generic scheduling policies. Through iterative simulations, the system produces a finalized assignment of speakers to slots, along with refined policies, updated email templates, and executable stop hooks that users can later deploy in real-world agentic workflows.
The system architecture comprises four key components: (1) a Simulation Engine that drives the iterative planning-execution loop through coordinated AI agents (Section B.1), (2) a Policy Module that governs agent reasoning and adapts to user preferences (Section B.2), (3) an Edge Case Detector that identifies situations beyond existing policy coverage and escalates them for user intervention (Section B.3), and (4) a Dashboard Visualization that provides real-time visual feedback on task state and progress (Section B.4).
As illustrated in Figure 4, these four components interact in a loop. The Policy Module supplies policies to the coordination agent inside the Simulation Engine, which in turn exchanges emails with simulated respondent agents to advance the scheduling task. When a respondent’s reply falls outside existing policy coverage, the Edge Case Detector escalates the issue to the user and feeds the resulting clarification back into the Policy Module as an updated policy. Throughout this cycle, the Dashboard Visualization renders the evolving task state so that the user can monitor progress and intervene at any point. We describe each component in detail below.
B.1. Simulation Engine
The Simulation Engine is the core of DoubleAgents, implementing the iterative planning-execution loop that drives the coordination process. It comprises three main subcomponents: the Coordination Agent, which reasons over the task and orchestrates actions; the Simulated Respondent Agents, which emulate human behavior to create a realistic environment; and the Context Management Module, which maintains the evolving problem state.
B.1.1. Coordination Agent
The coordination agent is the primary reasoning component of the system, responsible for assessing the current problem state, selecting relevant policies, generating plans, and executing actions. It operates within the ReAct framework and proceeds through the following stages at each simulation step:
  • Progress Understanding. The agent reviews the current problem context, including speaker information, email logs from previous days, and the history of user-approved plans and actions, and generates a concise summary of the current status. This summary indicates which days have been scheduled, what actions have been taken, and the most recent contact date for each speaker. The prompt used for generating the progress summary is provided in Appendix D.1.1.
  • Policy Selection. Based on the progress summary, the agent invokes the policy selection tool to identify the most relevant policies for the current state (Section B.2).
  • Plan Generation. Guided by the selected policies, the agent proposes a high-level plan (e.g., “Send email to all speakers to request availability for all seminar slots”). The plan and its rationale are presented to the user via the Interactive Interface (Figure 2(Middle)) for review and approval. Users may approve, edit, or request regeneration of the plan (Figure 7(a)).
  • Action Generation. Upon plan approval, the agent generates a set of concrete actions to operationalize the plan (Figure 7(b)). There are three types of actions: (a) drafting emails to selected respondents, (b) sending emails to specified recipients, and (c) awaiting responses from certain respondents. Users retain full control over actions and may edit, remove, or request regeneration. Detailed prompts for plan and action generation are provided in Appendix D.1.3.
  • Action Execution. The agent instantiates each approved action by invoking the appropriate tools from its toolbox (Figure 5). For email drafting, it calls the writeEmail() API, which generates emails given the current problem state, seminar details, user-provided templates, and historical email edits. Email drafting is performed using GPT-4o. For sending, it calls the sendEmail() API targeting the appropriate speaker email addresses. For delayed responses, it calls the waitResponse() tool, pausing execution and advancing to the next pending action or simulated day.
Users interact with the coordination agent through the Interactive Interface (Figure 2(Middle)), which displays the full history of plans, actions, and emails as interactive pop-up windows. The interface also supports natural language queries (e.g., “What did Prof. Lucy Tran reply in the last email?”), powered by GPT-4o (see Appendix D.4). We use GPT-o3 as the backbone of the coordination agent, given its strong reasoning capabilities.
B.1.2. Simulated Respondent Agents
Drawing inspiration from (Park et al., 2023), we employ LLMs to simulate each respondent agent, generating actions that align with their assigned personas. Upon receiving emails from the coordination agent, each simulated respondent first reasons over its persona and then generates an appropriate action. For instance, a respondent whose persona specifies “You are very busy and reply very late. Normally, it takes you three follow-up emails to respond” will delay its reply accordingly, while a respondent described as prompt in communication will reply immediately. If a respondent’s persona includes specific preferences (e.g., “prefers presenting over Zoom”), the generated response will reflect those preferences. Each respondent composes and sends email responses, which advance the scheduling process and enable the user to coordinate subsequent steps. The prompt for the speaker agent is detailed in Appendix D.3. We adopt GPT-4o for email response generation via the sendEmail() tool.
B.1.3. Context Management Module
The Context Management Module maintains all relevant contextual information throughout the scheduling process. DoubleAgents represents the problem context using JSON structures, where each key denotes a specific type of contextual information (e.g., speaker profiles, time constraints, email history) and each value contains the corresponding content. As shown in Figure 5, the problem context includes keys such as “time info” (reflecting seminar slots), speaker details, and communication logs. The context is dynamically updated at each step, incorporating user feedback, agent actions, and simulated responses to maintain an up-to-date representation of the task state.
B.2. Policy Module
The Policy Module serves as the central mechanism through which user values and preferences are encoded and operationalized within the system. Policies are natural language rules that guide the coordination agent’s reasoning and planning at each step of the scheduling process. They are displayed in the Policy Panel (Figure 2(Left)), where users can inspect, add, edit, or delete them at any time.
B.2.1. Seed Policies
To bootstrap the system with broadly applicable guidelines, the lead authors designed a set of six initial “seed policies” that address common scheduling challenges: (1) individuals being reluctant to share their full availability, (2) ambiguous or unclear responses, (3) the need for multiple follow-ups to elicit responses, (4) resolving scheduling conflicts between two speakers, (5) initiating the process by broadly reaching out to potential participants, and (6) determining when to cease follow-up attempts with non-responsive participants. Each policy is designed to address a specific challenge and serves as a default guideline that the coordination agent can draw upon when generating plans.
B.2.2. Policy Selection
At each planning step, the coordination agent invokes a policy selection tool that identifies the most relevant policies given the current problem state. For instance, at the beginning of the scheduling process when no speaker availabilities have been collected, the system selects the policy: “If we are just beginning and have not yet contacted any speakers for availability, then ask each speaker for their availability across all slots.” The selected policy is highlighted in the Policy Panel (Figure 2(A)), providing users with a transparent rationale for the agent’s subsequent plan. Details of the prompts used for policy selection are provided in the Appendix D.1.2.
B.2.3. Policy Refinement via Stop Hooks
Policies are not static; they evolve throughout the simulation process. When the Edge Case Detector (Section B.3) identifies a situation not covered by existing policies and the user provides a clarification, the system records this intervention as a stop hook and uses it to refine the policy set. For example, if a speaker requests to present via Zoom and the user specifies a preference for in-person attendance, this clarification is converted into a new policy explicitly stating that remote participation is not preferred. In this way, user interventions during edge cases are transformed into reusable system policies that persist across subsequent simulation steps and can be deployed in real-world workflows.
B.2.4. Implicit User Preference Learning
Beyond explicit policy refinement, the system also supports implicit learning of user preferences by observing user-initiated edits to system-generated plans, actions, or emails. The coordination agent summarizes each edit into a concise statement, logs it in a user edit history, and incorporates it into the problem context for subsequent planning steps. This enables the agent to evolve its behavior in alignment with user preferences over time without requiring explicit policy authoring. We leverage GPT-4o for the summarization process, with prompt details provided in Appendix D.1.5.
B.3. Edge Case Detector
The Edge Case Detector serves as a safeguard that identifies situations falling outside the coverage of existing policies and escalates them to the user for clarification. It is implemented as a tool that the coordination agent invokes to check whether an incoming message from a respondent agent can be addressed using the current policy set.
When the detector identifies an edge case—for example, a speaker requesting to present via Zoom when no existing policy addresses remote participation—it triggers a clarification request through the Interactive Chat Interface. The system presents the user with a concise description of the situation and a question prompting guidance (e.g., “Prof. Lucy Tran requests a Zoom option due to circumstances preventing in-person attendance. Should I confirm Zoom availability for their seminar or insist on in-person participation?”). The user then provides a natural language instruction (e.g., “Tell Prof. Lucy Tran we prefer in-person due to the potentially large audience”), which the coordination agent processes to generate an appropriate action.
Critically, each user clarification is recorded as a stop hook and used to refine the policy set (Section B.2). This mechanism ensures that (1) ambiguous or novel situations are never resolved autonomously without user input, (2) user interventions are preserved as reusable policies for future reference, and (3) the policy set grows iteratively to cover an expanding range of scenarios. The edge case detector is implemented with GPT-4o, with prompts detailed in Appendix D.2.
B.4. Dashboard Visualization
To support user understanding and foster alignment, DoubleAgents provides real-time visual feedback on the task state and system progress throughout the scheduling process. We implement a Flask-based dashboard interface designed to make the planning workflow interactive, legible, and actionable. The dashboard is composed of several layered visual components, shown in Figure 2 (Right):
  • Assignment Progress Tracker. This panel visualizes tentative and confirmed speaker-to-slot assignments. It highlights which seminar slots remain unfilled and which speakers have been successfully scheduled, enabling users to assess how close the current plan is to completion at a glance.
  • Calendar View. Anchored in simulated days, this temporal layer helps users reason about timing, deadlines, and follow-up actions. The calendar highlights both the current simulated day and the seminar dates, allowing prioritization of actions based on urgency and scheduling constraints.
  • Communication History. This panel displays the message history between the coordination agent and each speaker, offering a transparent view of the outreach and response status. It enables users to track ongoing communications, detect gaps, and assess follow-up needs.
These visual layers operate in parallel, allowing the system to efficiently track and update the state of each element in real time. At the end of each simulated day, the system generates a daily summary that highlights key events and decisions, using GPT-4o (with prompt details in Appendix D.5). The summary is presented within the Interactive Interface, providing users with a consolidated overview before advancing to the next simulated day.
Appendix C User Study Procedures and Data Analysis
@@
Appendix D DoubleAgents Prompts
D.1. Coordination Agent
D.1.1. Progress Summary
@@ D.1.4. Email Drafting
D.1.5. Implicit User Preference Learning
@@
D.2. Edge Case Detection
@@ D.3. Simulated Respondents
D.4. Free-form QA Chat
@@ D.5. Daily Summary
Appendix E Technical Evaluation Details
E.1. Policy Selection
@@ E.1.1. Experiment prompts
E.2. Edge Case Detection
E.2.1. Test suite construction
@@
E.2.2. Prompts to generate linguistic variations and test data
@@ E.2.3. Experiment prompts
Appendix F System Screenshots
(a) The pop-up window for a generated plan.
(b) The pop-up window for generated actions following the plan.
(c) The pop-up window for a generated email.
Figure 7. Examples for plan and action generation.
(a) Email sent by the simulated Prof. A.